50 g of water at 40 oC is placed in a well-insulated container. Neglecting heat losses, what mass of ice at 0 oC must be added to reduce the temperature of the water to 12 oC? The specific heat capacity of water is 4.2 × 10^3 J kg–1 K–1; the specific latent heat of fusion is 3.4 × 10^5 J kg–1.


1
Expert's answer
2021-03-29T09:00:08-0400
"Mc(T_w-T)=m(c(T_w-T)+R)\\
(50)(4200)(40-12)\\=(4200(12-0)+3.4\cdot10^5)m\\m=15\ g"